OpenStudy (snowcrystal):

Help please I will fan and medal Which radical expression is in simplified form?

OpenStudy (snowcrystal):

\[A.) \frac{ 11y }{ \sqrt{3}}\] \[B.) \frac{ \sqrt{6} }{ 5y }\] \[C.) \frac{ \sqrt{17} }{ \sqrt{4} }\] \[D.) \sqrt{\frac{ 25 }{ 81 }}\]

OpenStudy (snowcrystal):

@random231

random231 (random231):

wat do u think @SnowCrystal ?

OpenStudy (snowcrystal):

I am not sure at all I have never done this before

OpenStudy (anonymous):

I think it is B!!!(:

OpenStudy (snowcrystal):

thanks :) I just wish they made this stuff easier.